A softball diamond has the shape of a square wide sides 60 ft long. If 

a player is running from second base to third at a speed of 24 ft/sec, at 

what rate is his distance from home plate changing when she is 20 ft 

from the third ?


1
Expert's answer
2021-03-26T16:56:37-0400

Solution:


From base 1 to base 2 then to base 3 is a right-angled triangle.

When the runner is 20 feet from base 3, they are 40 feet from base 2

Letting the distance between the runner and the home plate be s we have:

s² = 60² + 40² = 5200

so,

"s=20\sqrt{13}" feet

Letting the distance from base 1 to base 2 be x, from base 2 to base 3 be y we have:

s² = x² + y²

so,

"2s\frac{ds}{dt}=2x\frac{dx}{dt}+2y\frac{dy}{dt}"


Now, with "s=20\sqrt{13}" , x = 60, "\frac{dx}{dt}=0" , "\frac{dy}{dt}=24" and y = 40 we have:


"(40\sqrt{13})\frac{ds}{dt}=2(60)(0)+2(40)(24)"


so, "(40\sqrt{13})\frac{ds}{dt}=" 1920 => "\frac{ds}{dt}=\frac{1920}{(40\sqrt{13})}"

or, "\frac{ds}{dt}=\frac{(48\sqrt{13})}{13}" => 13.3 "(\frac{ft}{sec})"